# MultiFinBen: A Multilingual, Multimodal, and Difficulty-Aware Benchmark for Financial LLM Evaluation

### v2

@@ Abstract
Recent advances in large language models (LLMs) have accelerated progress in financial NLP and applications, yet existing benchmarks remain limited to monolingual and unimodal settings, often over-relying on simple tasks and failing to reflect the complexity of real-world financial tasks. We introduce MultiFinBen, the first multilingual and multimodal benchmark tailored to the global financial domain, evaluating LLMs across modalities (text, vision, audio) and linguistic settings (monolingual, bilingual, multilingual) on domain-specific tasks. We introduce two novel tasks, including PolyFiQA-Easy and PolyFiQA-Expert, the first multilingual financial benchmarks requiring models to perform complex reasoning over mixed-language inputs; and EnglishOCR and SpanishOCR, the first OCR-embedded financial QA tasks challenging models to extract and reason over information from visual-text financial documents. Moreover, we propose a dynamic, difficulty-aware selection mechanism and curate a compact, balanced benchmark rather than simple aggregation existing datasets. Extensive evaluations of 22 state-of-the-art models reveal that even the strongest models, despite their general multimodal and multilingual capabilities, struggle dramatically when faced with complex cross-lingual and multimodal tasks in financial domain. MultiFinBen is publicly released to foster transparent, reproducible, and inclusive progress in financial studies and applications. [^1] Data: https://huggingface.co/datasets/TheFinAI/PolyFiQA-Easy
@@
Figure 1: Overview of MultiFinBen.
@@
BM EN ZH JA ES EL BI MU Vis OCR Au DA
Plutus peng2025plutus
DOCMATH zhao2023docmath
FLARE-ES zhang2024dolares
FinBen xie2024finbenholisticfinancialbenchmark
FinAudio cao2025finaudiobenchmarkaudiolarge
PIXIU xie2023pixiu
DOLFIN nakhlé2025dolfindocumentlevelfinancial
CFinBench nie2024cfinbenchcomprehensivechinesefinancial
MultiFinBen (ours)
@@ 1 Introduction
LLMs have demonstrated remarkable prowess in various tasks zhao2025mmvumeasuringexpertlevelmultidiscipline ; Chen_2025 , yet their applications to finance Liu2023FinGPT ; xie2023pixiu ; wu2023bloomberggpt require the continual push of the limits of model capabilities xiong2025flagtraderfusionllmagentgradientbased . Although there are already several specialized benchmarks xie2024finben ; li2024investorbenchbenchmarkfinancialdecisionmaking ; cao2025finaudiobenchmarkaudiolarge that evaluate LLM in core financial tasks, they still suffer from two critical limitations. First, they are overwhelmingly monolingual and monomodal. However, real-world financial applications involve mixed inputs that require multilingual semantic parsing, multimodal inputs of text, tables, charts, and audio, and cross-cultural contextual understanding. Secondly, existing benchmarks rely on simple aggregation without difficulty-aware selection, leading to over-weighting of easy and duplicated tasks. For example, in FinBen xie2024finben , 8 of 36 datasets are textual analysis tasks, with 7 being simple enough for zero-shot LLMs to exceed 60% accuracy, causing inflated overall scores and increased evaluation costs while failing to expose model weaknesses and providing meaningful insights.
To address these gaps, we introduce MultiFinBen, the first unified benchmark that spans three modalities (textual, visual, and audio), three linguistic settings (monolingual, bilingual, and multilingual), and seven task categories at three difficulty tiers, comprising 34 distinct datasets in five languages, including English, Chinese, Japanese, Spanish, and Greek. In contrast to benchmarks built on monolingual or cross-lingual corpora, we introduce the first multilingual financial task, comprising PolyFiQA-Easy and PolyFiQA-Expert, curated from authentic financial reports and news to require joint reasoning over mixed-language inputs, with questions designed and validated by financial analysts through multi-stage reviews to ensure quality and domain fidelity. For multimodal understanding, we develop the first OCR-embedded visual–text dataset covering balance-sheet tables, chart-laden slides, and complex financial imagery. Additionally, we propose a difficulty-aware selection to highlight frontier challenges and avoid overrepresentation by retaining one dataset per modality–language-task tier with the largest inter-model gap.
@@
We evaluate 22 frontier models on MultiFinBen. Overall, GPT-4o leads with 50.67%, while monomodal and monolingual models lag far behind. Unimodal models fail on unsupported inputs, underscoring the necessity of cross-modal reasoning. A 10.29 pp gap between multilingual (7.50%) and monolingual (17.79%) QA reveals that existing models still struggle with cross-lingual generalization in financial contexts. Our structured difficulty design exposes steep performance drops from 31.24% (easy) to 6.63% (hard), reflecting the gap between current model capabilities and real-world financial task complexity. Crucially, newly introduced datasets, such as PolyFiQA-Easy/Expert and the first financial OCR QA task, surface as among the hardest challenges, addressing modality and linguistic gaps overlooked by prior benchmarks. These results demonstrate that MultiFinBen not only reveals current model limitations in cross-modal, cross-lingual, and domain-specific applications but also provides a systematic framework to guide future model improvements and the development of harder and more realistic financial datasets.
2 MultiFinBen Benchmark
Table 2: Overview of MultiFinBen.
@@ 2 MultiFinBen Benchmark
Modality Language Task Dataset Source Size Metrics License Difficulty Text English IE SC mariko-etal-2020-financial Financial News 8,630 F1 CC BY 4.0 Easy IE FinRED sharma2022finred Webhose Financial News and Earning Call Transcrip 1,068 F1 Public Hard IE FINER-ORD shah2023finer Financial News 1,080 Entity F1 CC BY-NC 4.0 Medium TA Headlines sinha2021impact Financial News 2,283 AVE F1 CC BY-SA 3.0 Easy TA TSA cortis2017semeval Financial Microblog 561 Accuracy CC BY-NC-SA 4.0 Medium QA XBRL-Math chen2022convfinqa XBRL Agent 1,000 Accuracy CDLA-Permissive 2.0 Easy QA FinQA chen2021finqa Annual Reports 1,147 Accuracy MIT License Hard QA TATQA zhu2021tat Annual Reports 1,668 Accuracy MIT License Medium TG ECTSUM mukherjee2022ectsum Earnings Call Transcripts from the Motley Fool 495 ROUGE-1 Public Hard TG EDTSUM xie2023pixiu Financial News 2,000 ROUGE-1 Public Medium RM CCF feng2024empowering Credit Card Transactions 2,278 MCC DbCL v1.0 Medium FO BigData22 soun2022accurate Tweets, Historical Prices 1,470 MCC Public Medium DM MSFT yu2024finconsynthesizedllmmultiagent TradeTheEvent Dataset 222 SR Open Source Medium Chinese IE RRE jiajia-etal-2024-auditwen Regulatory Documents 117 Accuracy CC BY-NC Medium TA AIE jiajia-etal-2024-auditwen Regulatory Documents 1,573 Accuracy CC BY-NC Easy TA LNE jiajia-etal-2024-auditwen Regulatory Documents 218 Accuracy CC BY-NC Medium QA FinanceIQ 10.1145/3583780.3615285 Exam Questions 7,123 Accuracy CC BY-NC-SA 4.0 Medium Japanese TA chabsa kubo2018chabsa Securities Reports2 7,723 macro F1 CC BY 4.0 Hard Spanish TA MultiFin jorgensen2023multifin Article Headlines 368 Accuracy MIT License Medium TA TSA pan2023evaluation News Headlines 726 Accuracy Public Hard QA EFPA 6 Exam Questions 50 Accuracy Public Medium TG FNS-2023 7 Annual Reports 50 ROUGE-1 Public Hard Greek IE GRFinNUM peng2025plutus Annual Reports3 100 Entity F1 Public Medium TA GRMultiFin jorgensen-etal-2023-multifin Article Headlines 54 Accuracy CC BY-NC 4.0 Medium QA GRFinQA peng2025plutus Exam Questions 225 Accuracy Public Easy TG GRFNS-2023 10386228 Annual Reports 50 ROUGE-1 CC BY 4.0 Hard Bilingual TG DOLFIN nakhlé2025dolfindocumentlevelfinancial Fundinfo Financial Documents 1,932 Comet-da-22 MIT License Easy Mulfilingual QA PolyFiQA-Easy Financial Reports1 and News4 76 ROUGE-1 Public Hard TG PolyFiQA-Expert Financial Reports1 and News4 76 ROUGE-1 Public Hard Vision Spanish IE SpanishOCR Regulatory Documents5 13,845 ROUGE-1 Public Medium English IE EnglishOCR SEC EDGAR Company Filings1 7,961 ROUGE-1 Public Hard QA TableBench xie2024open SynthTabNet (Fintabnet and Marketing categories) 450 Accuracy Public Medium Audio English TG MDRM-test qin2019you Earnings Conference Calls 22,208 WER Public Medium TG FinAudioSum cao2025finaudiobenchmarkaudiolarge Earnings Conference Calls 64 ROUGE-L Public Hard

### v3

@@ Abstract
Real-world financial analysis involves information across multiple languages and modalities, from reports and news to scanned filings and meeting recordings. Yet most existing evaluations of LLMs in finance remain text-only, monolingual, and largely saturated by current models. To bridge these gaps, we present 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first expert-annotated multilingual (five languages) and multimodal (text, vision, audio) benchmark for evaluating LLMs in realistic financial contexts. 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: introduces two new task families: multilingual financial reasoning, which tests cross-lingual evidence integration from filings and news, and financial OCR, which extracts structured text from scanned documents containing tables and charts. Rather than aggregating all available datasets, we apply a structured, difficulty-aware selection based on advanced model performance, ensuring balanced challenge and removing redundant tasks. Evaluating 21 leading LLMs shows that even frontier multimodal models like GPT-4o achieve only 46.01% overall, stronger on vision and audio but dropping sharply in multilingual settings. These findings expose persistent limitations in multilingual, multimodal, and expert-level financial reasoning. All datasets, evaluation scripts, and leaderboards are publicly released.  [^1] Data: https://huggingface.co/datasets/TheFinAI/PolyFiQA-Easy
@@
0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:: Benchmarking Large Language Models for Multilingual and Multimodal Financial Application
Xueqing Peng1, Lingfei Qian1, Yan Wang1, Ruoyu Xiang6, Yueru He3, Yang Ren1, Mingyang Jiang1, Vincent Jim Zhang1, Yuqing Guo1, Jeff Zhao1, Huan He1, Yi Han4, Yun Feng17, Yuechen Jiang7, Yupeng Cao7, Haohang Li7, Yangyang Yu7, Xiaoyu Wang6, Penglei Gao8, Shengyuan Lin20, Keyi Wang3, Shanshan Yang7, Yilun Zhao2, Zhiwei Liu10, Peng Lu11, Jerry Huang11, Suyuchen Wang11, Triantafillos Papadopoulos15,18, Polydoros Giannouris10, Efstathia Soufleri18, Nuo Chen19, Zhiyang Deng7, Heming Fu1, Yijia Zhao1, Mingquan Lin13, Meikang Qiu14, Kaleb E. Smith5, Arman Cohan2, Xiao-Yang Liu3, Jimin Huang1, Guojun Xiong12, Alejandro Lopez-Lira9, Xi Chen6, Junichi Tsujii16, Jian-Yun Nie11, Sophia Ananiadou10,18, Qianqian Xie1,* 1The FinAI, 2Yale University, 3Columbia University, 4Georgia Institute of Technology, 5NVIDIA, 6New York University, 7Stevens Institute of Technology, 8Quantitative Health Sciences, Cleveland Clinic, 9University of Florida, 10University of Manchester, 11University of Montreal, 12Harvard University, 13University of Minnesota, 14Augusta University, 15Athens University of Economics and Business, 16National Institute of Advanced Industrial Science and Technology, 17Asian Development Bank, 18Archimedes, Athena Research Center, 19National University of Singapore, 20Carnegie Mellon University Correspondence: xqq.sincere@gmail.com
Figure 1: Overview of 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:.
Benchmark EN ZH JA ES EL BI MU Vis OCR Au DA Plutus Peng et al. (2025) CFinBench Nie et al. (2024) FLARE-ES Zhang et al. (2024b) FinBen Xie et al. (2024a) PIXIU Xie et al. (2023) DOLFIN Nakhlé et al. (2025) FinMME Luo et al. (2025) MME-Finance Gan et al. (2024) FinAudio Cao et al. (2025b) 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: (ours)
@@
Figure 2: Representation examples of PolyFiQA-Easy, EnglishOCR, and GreekOCR.
@@ 1 Introduction
LLMs have demonstrated remarkable prowess in various tasks Zhao et al. (2025); Chen et al. (2025b), yet their applications to finance Liu et al. (2023b); Xie et al. (2023); Wu et al. (2023) require the continual push of the limits of model capabilities Xiong et al. (2025). Although there are already several specialized benchmarks Xie et al. (2024b); Li et al. (2024); Cao et al. (2025b) that evaluate LLM in core financial tasks, they still suffer from three critical limitations:
  • Monolingual and monomodal scope. Prior benchmarks largely focus on English text only, whereas real-world financial applications demand cross-lingual understanding and multimodal reasoning across text, tables, charts, PDFs, and audio.
  • Imbalanced task difficulty. Existing benchmarks are typically constructed by simple aggregation without difficulty calibration. This leads to an overrepresentation of trivial tasks. For example, in FinBen Xie et al. (2024b), 7 out of 8 text classification datasets can already be solved above 60% accuracy in zero-shot, resulting in inflated scores and obscuring true model weaknesses.
  • Missing key financial scenarios. No prior benchmark addresses multilingual financial reasoning that integrates heterogeneous sources, nor OCR tasks that require extracting and reasoning over scanned financial PDFs. These scenarios are ubiquitous in practice yet entirely absent in prior benchmarks.
Benchmark Design.
To address these gaps, we introduce 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first unified benchmark that spans three modalities (text, vision, and audio), three linguistic settings (monolingual, bilingual, and multilingual), and seven task categories across three difficulty tiers. In total, it comprises 36 datasets in five languages (English, Chinese, Japanese, Spanish, and Greek). Unlike prior benchmarks limited to monolingual or unimodal corpora, 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: introduces two novel multilingual financial reasoning datasets, PolyFiQA-Easy and PolyFiQA-Expert, requiring joint reasoning over mixed-language inputs. In addition, we present the first financial OCR datasets (EnglishOCR, JapaneseOCR, SpanishOCR, and GreekOCR) to evaluate models on document-image understanding, a setting ubiquitous in practice but overlooked in prior work. Finally, we propose a difficulty-aware selection mechanism that retains one dataset per modality–language–task tier with the largest inter-model gap, avoiding redundancy and ensuring meaningful evaluation.
Evaluation and Findings.
@@
We benchmark 21 frontier models on 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:. Even the strongest model, GPT-4o, achieves only 46.01%, while monolingual and unimodal models fail dramatically on unsupported inputs, underscoring the necessity of cross-modal reasoning. Its significant low performance in multilingual scenario (9.79% and 5.31%) highlights persistent weaknesses in cross-lingual generalization. Our difficulty-aware design reveals steep drops from 32.73% (easy) to 7.20% (hard), exposing the gap between current model capabilities and real-world complexity. Crucially, our newly introduced datasets (PolyFiQA-Easy/Expert, EnglishOCR, and JapaneseOCR) emerge as among the hardest challenges, demonstrating how 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: systematically surfaces limitations and provides a roadmap for developing more realistic financial evaluation settings.
Our Contributions.
We summarize the main contributions of this work as follows:
$\triangleright$ First multilingual and multimodal benchmark for financial LLMs. 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: spans five languages (EN, ZH, JA, ES, EL), three modalities (text, vision, audio), and seven task categories, offering the most comprehensive evaluation setting.
$\triangleright$ Novel and realistic datasets. We introduce PolyFiQA-Easy and PolyFiQA-Expert, the first multilingual financial QA datasets requiring cross-lingual reasoning, as well as EnglishOCR, JapaneseOCR, SpanishOCR, and GreekOCR, the first financial OCR tasks targeting scanned financial documents.
$\triangleright$ Difficulty-aware evaluation framework. We propose a principled dataset selection mechanism based on inter-model performance gaps, ensuring balanced coverage across easy to hard tiers, in contrast to prior aggregation-based benchmarks.
$\triangleright$ Comprehensive model evaluation and insights. We evaluate 21 state-of-the-art models and reveal substantial performance gaps across modalities, languages, and difficulty levels, with even GPT-4o achieving only 46.01%. This demonstrates the necessity of our benchmark for stress-testing current models and guiding future development.
$\triangleright$ Open resources. We release all datasets and code to foster transparency, reproducibility, and future extensions by the research community.
2 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: Benchmark
In this section, we present 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first unified benchmark for evaluating LLMs in the financial domain across diverse modalities, linguistic settings, and tasks (Table  2).
@@
Modality Language Task Dataset Source Size Metric License Difficulty Text English IE SC Mariko et al. (2020) Financial News 8,630 F1 CC BY 4.0 Easy IE FinRED Sharma et al. (2022) Webhose Financial News and Earning Call Transcrip 1,068 F1 Public Hard IE FINER-ORD Shah et al. (2023) Financial News 1,080 Entity F1 CC BY-NC 4.0 Medium TA Headlines Sinha and Khandait (2021) Financial News 2,283 Avg F1 CC BY-SA 3.0 Easy TA TSA Cortis et al. (2017) Financial Microblog 561 Accuracy CC BY-NC-SA 4.0 Medium QA XBRL-Math Chen et al. (2022) XBRL Agent 1,000 Accuracy CDLA-Permissive 2.0 Easy QA FinQA Chen et al. (2021) Annual Reports 1,147 Accuracy MIT License Hard QA TATQA Zhu et al. (2021) Annual Reports 1,668 Accuracy MIT License Medium TG ECTSUM Mukherjee et al. (2022) Earnings Call Transcripts from the Motley Fool 495 ROUGE-1 Public Hard TG EDTSUM Xie et al. (2023) Financial News 2,000 ROUGE-1 Public Medium RM CCF Feng et al. (2024) Credit Card Transactions 2,278 MCC DbCL v1.0 Medium FO BigData22 Soun et al. (2022) Tweets, Historical Prices 1,470 MCC Public Medium DM MSFT Yu et al. (2024) TradeTheEvent Dataset 222 SR Open Source Medium Chinese IE RRE Huang et al. (2024a) Regulatory Documents 117 Accuracy CC BY-NC Medium TA AIE Huang et al. (2024a) Regulatory Documents 1,573 Accuracy CC BY-NC Easy TA LNE Huang et al. (2024a) Regulatory Documents 218 Accuracy CC BY-NC Medium QA FinanceIQ Zhang and Yang (2023) Exam Questions 7,123 Accuracy CC BY-NC-SA 4.0 Medium Japanese TA chabsa Kubo et al. (2018) Securities Reports 7,723 Macro F1 CC BY 4.0 Hard Spanish TA MultiFin Jørgensen et al. (2023a) Article Headlines 368 Accuracy MIT License Medium TA TSA Pan et al. (2023) News Headlines 726 Accuracy Public Hard QA EFPA 5 Exam Questions 50 Accuracy Public Medium TG FNS-2023 6 Annual Reports 50 ROUGE-1 Public Hard Greek IE GRFinNUM Peng et al. (2025) Annual Reports 100 Entity F1 Public Medium TA GRMultiFin Jørgensen et al. (2023b) Article Headlines 54 Accuracy CC BY-NC 4.0 Medium QA GRFinQA Peng et al. (2025) Exam Questions 225 Accuracy Public Easy TG GRFNS-2023 Zavitsanos et al. (2023) Annual Reports 50 ROUGE-1 CC BY 4.0 Hard Bilingual TG DOLFIN Nakhlé et al. (2025) Fundinfo Financial Documents 1,932 Comet-da-22 MIT License Easy Multilingual QA PolyFiQA-Easy Financial Reports1 and News7 172 ROUGE-1 Public Hard TG PolyFiQA-Expert Financial Reports1 and News7 172 ROUGE-1 Public Hard Vision English IE EnglishOCR SEC EDGAR Company Filings1 7,961 ROUGE-1 Public Hard QA TableBench Xie et al. (2024c) SynthTabNet (Fintabnet and Marketing categories) 450 Accuracy Public Medium Japanese IE JapaneseOCR FSA White Paper2 17,586 ROUGE-1 Public Hard Spanish IE SpanishOCR Regulatory Documents3 12,819 ROUGE-1 Public Medium Greek IE GreekOCR Annual Company Filings on Athens Stock Exchange4 6,533 ROUGE-1 Public Medium Audio English TG MDRM-test Qin and Yang (2019) Earnings Conference Calls 22,208 WER Public Medium TG FinAudioSum Cao et al. (2025b) Earnings Conference Calls 64 ROUGE-L Public Hard
Table 2: Overview of 0.44314 0.07059 0.66667M0.4902 0.06275 0.60392u0.54118 0.05882 0.5451l0.59216 0.05098 0.48235t0.64314 0.04314 0.42353i0.6902 0.03529 0.36078F0.74118 0.02745 0.30196i0.79216 0.01961 0.23922n0.84314 0.01569 0.18039B0.8902 0.00784 0.11765e0.94118 0 0.05882n\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:.